OpenStudy (anonymous):

What is sec(a + B) If cos(a+B) = cos a cos B - sin a sin B then does sec(a+ B) = 1 / cos(a+B) = 1 / cos a cosB - sin a sin B ??

OpenStudy (usukidoll):

I'm wondering if that exists... usually secant is 1/cosine... but I haven't seen a sec(a+b) unless I'm living under a rock.

OpenStudy (usukidoll):

that's sum and difference identities.

OpenStudy (anonymous):

Well, I know we have cos(a+B) = cos a cos B - sin a sin B which turns to \(cos(2\theta) \) and I have a problem I am working on that has \(sec(2\theta) \) so I was wondering

OpenStudy (anonymous):

or \(cos(2\theta) = cos^2-sin^2\)

zepdrix (zepdrix):

sec(a+ B) = 1 / cos(a+B) = 1 / (cos a cosB - sin a sin B) Yah this looks fine ^ But if you wanted to relate sec(A+B) to a bunch of secA and secB stuff, that would be a little bit more work :)

zepdrix (zepdrix):

What's the entire question? +_+ It's kind of hard to suggest which route you should take without seeing it

OpenStudy (anonymous):

This is the problem \( \sec(2\theta) = \frac{\csc^2\theta}{\csc^2\theta-2} \)

OpenStudy (anonymous):

This is the problem \( \Huge \sec(2\theta) = \frac{\csc^2\theta}{\csc^2\theta-2} \)

OpenStudy (anonymous):

I was thinking

OpenStudy (anonymous):

\[ \Huge \frac{1}{\cos^2\theta-\sin^2\theta} = \frac{\csc^2\theta}{\csc^2\theta-2} \] Going with it this way

zepdrix (zepdrix):

\[\Large\rm \sec(2\theta) = \frac{\csc^2\theta}{\csc^2\theta-2}\] \[\Large\rm \frac{1}{\cos(2\theta)} = \frac{\csc^2\theta}{\csc^2\theta-2}\] \[\Large\rm \frac{1}{1-2\sin^2 \theta} = \frac{\csc^2\theta}{\csc^2\theta-2}\]Recall that you can write the cosine double angle in like 3 different ways. I think THIS is the double angle identity that we want to use.

zepdrix (zepdrix):

See how the 2's match up? This one should get us on track.

zepdrix (zepdrix):

The next step is a little tricky ;) Ok with that step though? :O

OpenStudy (anonymous):

Yes! I forgot about that one. Do we \( \Large\rm \frac{1}{1-2\sin^2 \theta}\times\frac{1+2\sin^2 \theta}{1+2\sin^2 \theta} \)

zepdrix (zepdrix):

Conjugate? Hmm no that's going to give us a 4 in the bottom, i don't think we want that

zepdrix (zepdrix):

What is csc(theta) in relation to sin(theta)? \(\Large\rm \csc(\theta)=\frac{1}{\sin(\theta)}\), ya? Hmm so how can we `end up with` a 1/sin^2x in the top there? :d

OpenStudy (anonymous):

Multiply by its reciprocal

zepdrix (zepdrix):

Well if want to turn a 1 into 1/sin^2 then we should probably divide by sin^2. So let's divide top `and` bottom by sin^2(theta). I know it's a weird step!

zepdrix (zepdrix):

\[\Large\rm \frac{1}{1-2\sin^2 \theta} \color{orangered}{\left(\frac{\frac{1}{\sin^2\theta}}{\frac{1}{\sin^2\theta}}\right)}= \frac{\csc^2\theta}{\csc^2\theta-2}\]

OpenStudy (anonymous):

Yes that is what I got and you are correct. That is weird :-D

zepdrix (zepdrix):

It probably would have made more sense to start with the right side and turn it into the left. The weird step in the middle would had been more straight forward :) This is fine though

OpenStudy (anonymous):

Cool, I got it. Thank you!!!

OpenStudy (anonymous):

Working it out the rest of the way

zepdrix (zepdrix):

cool :)

OpenStudy (anonymous):

I think the ones with the sec and csc gives me the most trouble